using the theorem what is the answer to a problem where triangle sides are 3.1, 8.7 and x

Answer by checkley75(3666)

IT DEPENDS WEITHER THE X IS THE HYPOTENUSE OR ONE OF THE SIDES OF A RIGHT TRIANGLE.
IF X IS THE HYPOTENUSE THEN:
3.1^2+8.7^2=X^2
9.61+75.69=X^2
85.3=X^2
X=SQRT85.3
X=9.236 ANSWER.
IF X IS ONE OF THE SIDES THEN:
X^2+3.1^2=8.7^2
X^2+9.61=75.69
X^2=75.69-9.61
X^2=66.08
X=SQRT66.08
X=8.129 ANSWER.